Janet invested $13,000, part at 13% and part at 6%. If the total interest at the end of the year is $1,200, how much did she invest at each rate?
Answer by marcsam823(57)

Let x = total amount invested at 13%
Let y = total amount invested at 6%
x + y = 13000
y = 13000 - x
.13x = total amount of interest at 13%
.06y = total amount of interest at 6%
.13x + .06y = 1200
Substitute:
.13x + .06(13000 - x)= 1200
Multiply both sides by 100
13x + 6 (13000 - x) = 120000
Solve for x:
13x + 48000 - 6x = 120000
7x = 72000
x = 10285.71
Solve for y:
13000 - x = y = 2714.28
Janet invested 10,285.71 at 13% and 2,714.28 at 6%